OpenStudy (anonymous):

Find a cubic function f(x)=ax^3+cx^2+d that has a local maximum value of 10 at -3 and a local minimum value of 6 at 0 what are a b c

OpenStudy (ash2326):

We have \[f(x)=ax^3+cx^2+d\] It has local maximum of 10 at x=-3 and local minimum of 6 at 0 so \[10=-27a+9c+d\] and \[6=d\] Wait a minute Where is b in f(x)???

OpenStudy (experimentx):

f(x)=ax^3+cx^2+d f'(x) = 3ax^2+2cx since it has minima at -3 and maxima at 0 f'(-3) = 0 f(-3) = 10 f'(0) = 0 f(0) = 6 3 variables 4 equations, should be solvable.